$SPAD/src/input bbtree.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool bbtree.output
)set message test on
)set message auto off
)clear all
--S 1 of 10
lm := [3,5,7,11]
--R
--R
--R   (1)  [3,5,7,11]
--R                                                  Type: List(PositiveInteger)
--E 1
--S 2 of 10
modTree(12,lm)
--R
--R
--R   (2)  [0,2,5,1]
--R                                                          Type: List(Integer)
--E 2
--S 3 of 10
t := balancedBinaryTree(#lm, 0)
--R
--R
--R   (3)  [[0,0,0],0,[0,0,0]]
--R                                 Type: BalancedBinaryTree(NonNegativeInteger)
--E 3
--S 4 of 10
setleaves!(t,lm)
--R
--R
--R   (4)  [[3,0,5],0,[7,0,11]]
--R                                 Type: BalancedBinaryTree(NonNegativeInteger)
--E 4
--S 5 of 10
mapUp!(t,_*)

--R
--R
--R   (5)  1155
--R                                                        Type: PositiveInteger
--E 5
--S 6 of 10
t
--R
--R
--R   (6)  [[3,15,5],1155,[7,77,11]]
--R                                 Type: BalancedBinaryTree(NonNegativeInteger)
--E 6
--S 7 of 10
mapDown!(t,12,_rem)
--R
--R
--R   (7)  [[0,12,2],12,[5,12,1]]
--R                                 Type: BalancedBinaryTree(NonNegativeInteger)
--E 7
--S 8 of 10
leaves %
--R
--R
--R   (8)  [0,2,5,1]
--R                                               Type: List(NonNegativeInteger)
--E 8
--S 9 of 10
squares := [x**2 rem m for x in % for m in lm]
--R
--R
--R   (9)  [0,4,4,1]
--R                                               Type: List(NonNegativeInteger)
--E 9
--S 10 of 10
chineseRemainder(%,lm)
--R
--R
--R   (10)  144
--R                                                        Type: PositiveInteger
--E 10
)spool
)lisp (bye)

———-

References^{[1] nothing}

5